Journalist Report – February 24th

Sol 0

Author : Benjamin Auzou, Journalist

Title : "Our new Home"

The six of us woke up after a difficult night in the Science Dome,
because of the lack of comfort but above all because of our excitation
to begin our mission. The whole crew chose to begin the day with a
sport session ! That was a difficult session for some of us, but we
all needed that to wake up and prepare the three weeks of mission
ahead of us.

At 0800 the crew 205 left the Hab to return to their five countries.
Natalia, Dave, Nathan, Maria, Ghenim, Daniel, Hannah and Veronica left
us the station after a full afternoon of training yesterday. We wish
you a good continuation ! You’ve done a good job here in the MDRS,
even if the martian weather wasn’t clement with you. We will continue
to hold the station well and make great science here on Mars !

Then we spend the whole morning taking our marks in the Hab, we took
our first breakfast composed of milk and scrambled eggs. We chose our
rooms, our tiny zones of privacy in the Hab.

For the rest of the day, waiting for Atila to train us, we all worked
on our experiments or on the Hab’s organization :
Norbert, our GreenHab Officer, spend a lot of time in the GreenHab to
discover the plants in here and understand their needs. Cerise, our
crew Biologist and Health and Safety Officer tested everything in the
Campus to assure our safety : alarms, fire extinguishers, medical
supplies and radios. Gaspard, our engineer and the "hacker" of the
team tackled the internet consumption of our devices to avoid data
losses (we only used 100MB up to this time!). In the same time,
Aurélien and I organized in the kitchen the food supplies that arrived
with us. And Jérémy, Commander of our mission prepared all the
material necessary to follow our water consumption.

At 1430 pm, Atila arrived to train us with the rovers and gave us his
last advice.

We had a long discussion on our goals for the mission. And that was
the moment to close the airlock door and begin our simulation.

I am Benjamin, journalist of the crew 206, million of kilometers away
from Earth, in the Mars Desert Research Station. The weather here is
beautiful, even if the soil is still damaged by the storm that stroke
the crew 205’s mission. My crew and I will sleep well after this
tiring day. I am in a hurry to wake up tomorrow to start experiments
and Extravehicular Activities, to live the Martian Dream.

Operations Report – February 24th

SOL: 0

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need
replacement on generator. A contractor has been arranged to make these
repairs within a few days.

Generator (hours run): Turned off at 1430 (24th Feb), Turned on at
1903 (24th Feb)

Solar— SOC % – Turned off Gen (24th Feb) 75%, Turned on Gen (24th Feb) 100%

Diesel Reading – ~ 75%

Propane Reading – ~ 26%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – around 76%, ~ 419 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 152,9 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01415558 units

Toilet tank emptied: Yes

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Used

Hours: 51.3 hours

Beginning charge: 100%

Ending charge: 74%

Currently charging: Yes

Curiosity rover used: Used

Hours: 75.5 hours

Beginning charge: 100%

Ending charge: 75.5%

Currently charging: Yes

Spirit rover used: Used

Hours: 78.3 hours

Beginning charge: 100%

Ending charge: 66%

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Used to bring back crew 205 to Grand Junction

Crew Car used and why, where? Used to bring back crew 205 to Grand Junction

General notes and comments: In the Hab’s upper deck, the bulb from one
of the lamps (on the sink side) is menacing to fall, one screw is not
holding anymore. Also this lamp is holding onto the electrical cable
rather than its theoretical link to the ceiling.

Summary of internet: Nothing to report

Summary of suits and radios: New inventory of the radios this morning, five of them have minor to major issues. Going to test them again tomorrow after full charging.

Summary of Hab operations: Nothing to report

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Propane tank
needs to be filled rapidly, how do we proceed ?

GreenHab Report – February 24th

GreenHab Officer: Norbert Pouzin

Environmental control: Heating and Cooling w/ ambient air (0 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 17.0°C

High: 33.2°C

Hours of supplemental light: 5 Hours

Daily water usage for crops: 6.6 gallons

Water in Blue Tank: ~ 152.9 gallons

Crops watered at 0905 hours (3.3 gallons) and 1800 hours (3.3 gallons):


1) 09.05 AM 21.8° C / 18 %
2) 12.00 AM 33.0° C / 0 %
3) 14:30 PM 19.7° C / 0 %
4) 18:00 PM 21.9° C / 30 %

Changes to crops: None.

Narrative: The door was opened from 1200 to 1700 to cool down the
greenhab, since the temperature reached 33°C.
After our formation with Atila, and with his agreement, we decided to
move the grow tent from the science dome to the greenhab, so I could
start my plant experiment tomorrow (study of the influence of music,
to improve radish defence mechanism to radiation).
A plant stake was added for the chives.
The blue tank wasn’t fil today, it will be tomorrow.
7.0g of basil were harvested for soup.

Harvest: 7.0 g basil

Support/supplies needed: None.

Sol Summary – February 24th

Sol: 0

Summary Title: First Marks

Sol Activity Summary: After a small and difficult night of sleep in
the science dome (we let the leaving crew in the hab so they could
finish packing their bags), we got out of "bed" at 7, to say hi to
crew 205 and to do our first daily physical training. We ate breakfast
before unpacking our bags and getting settled in the hab. At 10 we
started preparing our experiments (e.g. water consumption monitoring)
and working on the food inventory.

We prepared lunch around 12h30 and ate until 2, when Atila came for our
training. The training went fine, we learned the last few details for
the mission. After that we each continued working on our experiments
and our reports, to be on time for the COMMS.

Anomalies in work: none

Weather: sunny

Crew Physical status: fine

Copyright © The Mars Society. All rights reserved. | Main Site